During 2021, a company sells 160 units of inventory for $55 each. The company has the following inventory purchase transactions for 2021:

Date Transaction Number
of Units
Unit
Cost
Total Cost
Jan. 1 Beginning inventory 58 $ 40 $ 2,320
May 5 Purchase 107 39 4,173
Nov. 3 Purchase 75 38 2,850
240 $ 9,343


Actual sales by the company include its entire beginning inventory, 87 units of inventory from the May 5 purchase, and 15 units from the November 3 purchase.


Calculate cost of goods sold and ending inventory for 2021 assuming the company uses specific identification.


Cost of goods sold:

Ending inventory:
Answer #1
Cost of goods sold:
From beginning inventory 2320 =58*40
From May 5 purchase 3393 =87*39
From November 3 purchase 570 =15*38
Cost of goods sold 6283
Ending inventory:
From beginning inventory 0
From May 5 purchase 780 =(107-87)*39
From November 3 purchase 2280 =(75-15)*38
Ending inventory 3060